OpenStudy (anonymous):

What is the sum of a 11–term geometric series if the first term is 4 and the last term is 4,096?

OpenStudy (amistre64):

\[a_n=4\frac{1-r^n}{1-r}\] i believe is the general setup for this

OpenStudy (amistre64):

or that might be the sum set up :/

OpenStudy (callisto):

\[r=\sqrt[10]{4096/4} = 2\] use the formula above and solve it :)

OpenStudy (amistre64):

\[a_n=4*r^{n-1};\ n=1,2,3,...\]

OpenStudy (callisto):

wait, it should be (r^n -1)/(r-1)

OpenStudy (amistre64):

same, you only take a negative out of the top and tranfer it to the bottom or visversa

OpenStudy (callisto):

not really lol take the limit and you'll see (i made the same mistake before lol)

OpenStudy (anonymous):

ok so what goes where. thats my problemm

OpenStudy (amistre64):

\[\frac{1-r^n}{1-r}*\frac{-1}{-1}=\frac{r^n-1}{r-1}\]

OpenStudy (anonymous):

i think you need to find r first maybe?

OpenStudy (callisto):

sum = a (r^n -1) / (r-1) where r = 2

OpenStudy (callisto):

a=4

OpenStudy (anonymous):

\[4, 4r, 4r^2, ..., 4r^{10}\] so you have \[4r^{10}=4,096\] \[r^{10}=1024\] \[r=\sqrt[10]{1024}=2\]

OpenStudy (anonymous):

so r is 2?

OpenStudy (anonymous):

oops callisto wrote that already yes r = 2

OpenStudy (amistre64):

i like how they ask: is that the results? after youve written out how you get the results ... typoes persist tho

OpenStudy (callisto):

alright answer = 8188

OpenStudy (anonymous):

\[4\times \frac{2^{11}-1}{2-1}=4(2^{11}-1)\] is your sum

OpenStudy (anonymous):

what callisto said

OpenStudy (callisto):

amistre64 when you take the limit it's really different!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

8184

OpenStudy (callisto):

8188!!!

OpenStudy (anonymous):

ok well i was close

OpenStudy (amistre64):

@Callisto the wolfs being obstinate http://www.wolframalpha.com/input/?i=limit%28n+to+inf%29+%281-r%5En%29%2F%281-r%29

OpenStudy (callisto):

the first post is the mistake i made, the second post is the limit part